OPTICAL TOMOGRAPHY IN TWO DIMENSIONS

PLAMEN STEFANOV GUNTHER UHLMANN

Abstract. We consider in two dimensions, the inverse boundary problem of reconstructing the absorption and scattering coefficient of an inhomogeneous medium by probing it with diffuse light. The problem is modeled as an inverse boundary problem for the stationary linear Boltzmann equation. The information is encoded in the albedo operator. We show that we can recover the absorption and the scattering kernel from this information provided that the latter is small in an appropriate topology. We also give stability estimates.